OpenStudy (anonymous):

can someone check my answer What is the solution to the rational equation x/(x^2-9)+1/(x+3)=1/(4x-12) I got -9/7 for my answer am i correct???

OpenStudy (anonymous):

@Angels

OpenStudy (anonymous):

x= 15/7 that wut i got

OpenStudy (anonymous):

thanks that was my second choice

OpenStudy (anonymous):

ur welcome

OpenStudy (phi):

\[ \frac{x}{x^2-9} + \frac{1}{x+3} = \frac{1}{4x-12} \] it looks like you should factor things, \[ \frac{x}{(x-3)(x+3)} + \frac{1}{x+3} = \frac{1}{4(x-3)} \] we could find a common denominator, or we could multiply to "clear the denominators" for example, if we multiply by the common denominator 4(x-3)(x+3) we get \[ \frac{4x(x-3)(x+3)}{(x-3)(x+3)} + \frac{4(x-3)(x+3)}{x+3} = \frac{4(x-3)(x+3)}{4(x-3)} \\ \frac{4x\cancel{(x-3)(x+3)}}{\cancel{(x-3)(x+3)}} + \frac{4(x-3)\cancel{(x+3)}}{\cancel{(x+3)}} = \frac{\cancel{4(x-3)}(x+3)}{\cancel{4(x-3)}} \] that turns into \[ 4x + 4(x-3) = x+3\\ 4x+4x-12 = x+3\\ 7x= 15\\x= \frac{15}{7}\]